OpenStudy (anonymous):

A toaster is plugged into an 120-volt outlet. If it has a resistance of 460 ohms, how much power does it use? 1,800 watts 3.8 watts 31 watts 55,000 watts 340 watts

OpenStudy (anonymous):

do you u have a formula?

OpenStudy (anonymous):

@rg40016

OpenStudy (anonymous):

no :/ I got it off from plato

OpenStudy (anonymous):

hmm let me see if I can find one.

OpenStudy (anonymous):

use this formula: power = amps (a) voltage = (v) ohms = (o) voltage / ohms = amps or power insert the known values and solve.

OpenStudy (kropot72):

The simplest way to find the power is to use the formula: \[\large P=\frac{V^{2}}{R}=\frac{120^{2}}{460}\ watts\]